• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Feuerbach

    Log entry time 17:12:15 on June 09, 2005

    Entry number 145629

    keyword=Plan for 9-June through 10-June

    Plan for 9-June through 10-June

    The RICH is almost installed, and will need to be flushed before it is
    used in "production mode."

    SPECIAL NOTES:
    - if there is ANY problem with the L-HRS magnet controls, BE CERTAIN to
    Halog it. The Techs have upgraded the controller and believe it should be
    fixed, so record any contrary evidence to permit debugging.

    - Bodo and other RICH experts have free-reign to do what needs to be
    done to check the status/performance of the RICH. The HV is NOT to be
    turned on until they say so.

    TASKS:
    ------
    1) Task: Compare SLI and OTR (1C12) measurements of the beam-energy
    spread. Keep in mind that we are at the edge of the OTR's resolution, so
    it will be acting as an UPPER BOUND on the energy spread.
    a) Once beam is stable, inform MCC that we will be wanting to measure
    the beam-energy spread for different currents.
    b) Request beam current of 25 microAmps; request the OTR at 1C12 to be
    inserted -- make sure this doesn't affect the SLI result.
    c) When beam is stable, take a SIMULTANEOUS snapshot of the 1C12 OTR
    and SLI screens, and immediately post to the HALOG, noting the current.
    d) Repeat b) and c) for 50, 75, and 100 microamp running.

    2)"Production" running on the thick (100mg/cm^2) C-12 target. Alternate
    between taking data at 50 and 100 microamps, two hours at each setting.
    a) When switching currents, have the 1C12 OTR inserted and record the
    OTR and SLI screens.
    b) For each run, do a detector replay (50000 events) and a physics
    replay. See Notes below.
    c) The energy-spread from the SLI SHOULD BE 3.e-5 or LESS. If it gets
    to be 4.e-5, ask MCC to improve it. If it cannot be improved,
    contact the Run Coordinator.


    Production Running Notes:
    -------------------------

    Set the spectrometers to the production momenta (see the whiteboard); the
    septum magnets should be automatically set to the proper set-current.

    1) Analyze the data looking to see the RF buckets in coincidence timing.
    Notes are in /adaqfs/home/adaq/e94107/onlana/analysis_hints.html

    2) We are starting with the prescales from run 2200 in January of 2004.
    This running used 50 uA on the same waterfall target and should be a good
    starting point. Adjust these prescales at 50 uA to get: All the T7
    triggers, as many T5 triggers as possible while keeping total event
    recording rate below 0.6 KHz, about 10 Hz each of T1, T3, T6, and 5 Hz of
    T2 and T4. HALOG the prescales. Remember this will change to about 1.2
    KHz at max current.

    3) Make hourly HALOGs of the SLI status and stripline chart of energy
    spread. In shift summary, keep rough track of frequency of beam trips.
    HALOG the BCM showing the 8-hours of the shift.


    When in doubt, contact Rob Feuerbach (RC cell phone: 876-1787; page
    584-5111).